# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

### v1

@@ VI Conclusion and future works
In this paper, we propose a CNS-inspired control framework based on SNN, enabling agile robotic arm control in complex environments. The framework introduces a hierarchical architecture with five modules, is fully implemented using SNNs, and has been validated through both simulations and real-world robotic experiments, outperforming an industrial-grade baseline. Future work will focus on improving reinforcement learning for higher reproducibility, enhancing the cerebellar module for better gravity compensation, and further increasing control accuracy.
@@
[Uncaptioned image] Yanbo Pang received the B.Eng. from Sun Yat-sen University in 2022 and the M.Sc. from Tsinghua University in 2025. He is currently a Ph.D student at the Technical University of Munich (TUM). His research interests include neuromorphic computing, neurorobotics and robotic arm control.
[…]

### v2

@@ VI Conclusion and future works
In this paper, we propose a CNS-inspired control framework based on SNN, enabling agile robotic arm control in complex environments. The framework introduces a hierarchical architecture with five modules implemented using SNNs, and has been validated through both simulations and real-world robotic experiments. Results show that this framework operates in a manner similar to the human CNS and can outperform an industrial-grade baseline. Future work will focus on improving reinforcement learning for higher reproducibility, enhancing the cerebellar module for better gravity compensation, and further increasing control accuracy.
@@
Yanbo Pang received the B.Eng. from Sun Yat-sen University in 2022 and the M.Sc. from Tsinghua University in 2025. He is currently an intern at Booster Robotics. His research interests include neuromorphic computing, neurorobotics and robotic arm control.
Qingkai Li received his B.S. from Tsinghua University in 2021. He is currently a Ph.D. candidate in the Department of Automation at Tsinghua University, with research interests spanning neuromorphic computing, bio-inspired algorithms, imitation learning, and whole-body control.
Mingguo Zhao received the B.Eng., M.Sc. and Ph.D. from Harbin Institute of Technology in 1995, 1997 and 2001, respectively. He is currently a Full Professor with the Department of Automation and a director of the Beijing Innovation Center for Future Chips, Tsinghua University. His research interests include brain-inspired robot control and general robotic platform based on neuromorphic computing and hardware.